# AerialBooth: Mutual Information Guidance for Text Controlled Aerial View Synthesis from a Single Image

### v1

@@ 4 Experiments and Results
Figure 3: We show comparisons with the state-of-the-art image editing, personalization and text-controlled aerial view synthesis methods on the AerialBooth-Syn dataset. Our methods is able to generate images that are “more aerial”, while preserving the fidelity w.r.t. the input image.
Figure 4: We show comparisons with the state-of-the-art image editing, personalization and text-controlled aerial view synthesis methods on the AerialBooth-Real dataset. Our method achieves the best trade-off between generating images that are aerial and preserving the fidelity w.r.t. the input image.
Method CLIP A-CLIP SSCD DINO CLIP-I CLIPD A-CLIPD
Dataset: AerialBooth - Syn
DreamBooth LoRA $0.3069$ $0.1765$ $0.6260$ $0.6204$ $0.9058$ $0.0369$ $0.0334$
Imagic $0.3046$ $0.1748$ $0.5559$ $0.5846$ $0.9064$ $0.0315$ $0.0329$
Aerial Diffusion $0.2920$ $0.2082$ $0.2056$ $0.2946$ $0.751$ $0.0674$ $0.0905$
Ours ($\lambda_{MI}=1e-5$) $0.3107$ $0.1859$ $0.3921$ $0.4493$ $0.8636$ $0.0512$ $0.0486$
Ours ($\lambda_{MI}=1e-6$) $0.3108$ $0.1862$ $0.3928$ $0.4532$ $0.8678$ $0.0520$ $0.0469$
Dataset: AerialBooth - Real
DreamBooth LoRA $0.2929$ $0.1791$ $0.5501$ $0.5088$ $0.9000$ $0.0371$ $0.0332$
Imagic $0.2960$ $0.1773$ $0.5024$ $0.5048$ $0.9065$ $0.0340$ $0.0348$
Aerial Diffusion $0.2896$ $0.2084$ $0.1717$ $0.2614$ $0.7606$ $0.0711$ $0.0721$
Ours ($\lambda_{MI}=1e-5$) $0.3047$ $0.1887$ $0.3326$ $0.3884$ $0.8700$ $0.0547$ $0.0453$
Ours ($\lambda_{MI}=1e-6$) $0.3045$ $0.1871$ $0.3345$ $0.3915$ $0.8742$ $0.0564$ $0.0429$
Table 1: We compare AerialBooth with various state-of-the-art methods on quantitative metrics indicative of viewpoint, fidelity and directional similarity. Our method achieves the best viewpoint-fidelity trade-off amongst all prior work.
In this section, we present the results of our method along with extensive comparisons and ablations.
@@
We collect a synthetic dataset, AerialBooth-Syn and a real dataset, AerialBooth-Real. Both datasets contain images across a wide variety of categories including indoor scenes, natural scenes, human actions, birds/ animals, animations, traffic scenes and architectures. AerialBooth-Syn contains $500$ images that were generated using Stable Diffusion XL [41]. To generate the text prompts for the generated images in AerialBooth-Syn, we used Large Languuage Models (LLMs) such as ChatGPT and Bard. AerialBooth-Real contains $139$ images downloaded from the Unsplash website, the text descriptions for these images were obtained using the BLIP-2 [29] model.
@@
We use the stable diffusion 2.1 model in all our experiments, ablations and comparisons. All our images (except for images in AerialBooth-Real) are at a resolution of $512\times 512$. With respect to $I_{S}$, we train the text embedding and the diffusion model for $1,000$ and $500$ iterations respectively, at a learning rate of $31e-$ 3 and $42e-$ 4 respectively. Using $1000$ iterations to optimize the text embedding ensures that the text embedding $e_{opt}$ at which $I_{S}$ is reconstructed is not too close to $e_{S}$, which would make it biased towards $I_{S}$ otherwise. Similarly, it is not too far from $e_{S}$ either, hence the text embedding space learns the characteristics of $I_{S}$. With respect to $I_{H}$, we train the text embedding and the diffusion UNet for $500$ and $250$ iterations respectively. We want $e_{H}$ to be in close vicinity of $e_{opt}$; we train the diffusion model for just $250$ iterations so that the model does not completely overfit to $I_{H}$. The role of $I_{H}$ is to create variance and provide pseudo supervision, it is not an accurate approximation of the aerial view. We set the hyperparameter for mutual information guidance at $51e-$ 5 or $61e-$ 6, the inference is run for $50$ steps.
@@
No concrete evaluation metrics exist for this task. We follow prior work on text-based image editing and personalization to evaluate our method on the following $7$ metrics:
  • Viewpoint and text alignment: We use the text description ‘aerial view, ’ + $t_{S}$ and ‘aerial view’, along with the generated image, to compute the CLIP-Score [44] and the A-CLIP Score respectively. The former indicates alignment of the generated image with detailed textual description of the image; the latter focuses more on the viewpoint.
  • Fidelity: To evaluate the alignment of the contents of the generated aerial-view image with the input image, we compute the CLIP-I score [50] which measures the cosine similarity between the embeddings of the aerial-view image and the input image in the CLIP space. We also use the self-supervised similarity detection metrics, DINO [6] and SSCD [40].
  • CLIP Directional Similarity: CLIP directional similarity [14] measures the consistency of the change between two images, $I_{S}$ and $I_{T}$, in the CLIP space with the change between the two image captions (dictating the transformation from $I_{S}$ to $I_{T}$). We compute two versions of this score, CLIPD score and the A-CLIPD score. CLIPD score uses ‘aerial view, ’ + text as the target text, and ‘A-CLIPD score’ uses ‘aerial view’ as the target text.
Higher values are desired for each of these metrics. Viewpoint faithfulness and fidelity w.r.t input image are a direct result of the bias-variance trade-off of the model. High values for both are desired. However, as noted by Blau et. al. [4], maximizing both is not straightforward; inevitably one of the factors will degrade in response to the improvement in the other. We generate $5$ aerial images, with random noise initializations, and choose the image with the highest CLIP + SSCD score (since CLIP is an indicator of viewpoint alignment and SSCD score measures the fidelity w.r.t. the input image).
Figure 5: We conduct ablation experiments to show the effectiveness of Inverse Perspective Mapping (IPM) and the mutual information guidance. IPM helps in the generation of images that are aerial, and mutual information guidance better preserves the contents of the input image.
State-of-the-art comparisons.
@@
4.1 Results
In this subsection, we analyze the results of our method and prior art on AerialBooth-Syn and AerialBooth-Real.
@@ 4.1 Results
We show qualitative results on AerialBooth-Syn in the first row of Figure 1 and in Figure 3. We show qualitative results on AerialBooth-Real in the second row of Figure 1 and in Figure 4. Our method is able to generate aerial views of the input images across a diverse set of scenes. Our method generates results that are more aerial in viewpoint than DreamBooth and Imagic, while being consistent with the contents of the input image. Aerial Diffusion is unable to generate good quality images of scenes that have many objects. Our method is able to deal with complex scenes as well as modify the viewpoint.
@@
We show the quantitative results in Table 1. Our method achieves a higher CLIP score than all prior work, indicating that it is able to generate an aerial view of the scene with contents dictated by the text better than prior work. The A-CLIP score achieved by our method is higher than that of DreamBooth and Imagic, indicating better conformance to the aerial viewpoint. Even though the A-CLIP score of AerialBooth is lower than than of Aerial Diffusion, Aerial Diffusion generates poor quality images for scenes with more than one object and also has diagonal artifacts in its generated images, as we observe from the other metrics and qualitative results, thus, offsetting its high A-CLIP Score.
@@
CLIP-I and self-supervised metrics such as SSCD and DINO are not viewpoint invariant. SSCD is a better metric [40] than DINO. In many cases, since Imagic and DreamBooth generate views close to the input view, instead of an aerial views, it is natural for them to have higher SSCD and DINO Score. Our method has a much higher SSCD and DINO score than Aerial Diffusion, showing considerable improvement over prior work in retaining the fidelity w.r.t. the input image, while modifying the viewpoint.
CLIPD and A-CLIPD measure the directional change in the generated image w.r.t. the text prompt describing the transformation. Higher values indicate the image moving more towards the aerial viewpoint. Our method achieves higher CLIP scores than Imagic and DreamBooth, indicating its success in generating aerial images. Synonymous to our analysis for CLIP and A-CLIP, while the CLIPD and A-CLIPD scores are higher for Aerial Diffusion than our method, Aerial Diffusion is unable to generate good quality images as evidenced by the CLIP-I score.
In summary, AerialBooth achieves the best viewpoint-fidelity trade-off amongst all prior work.
@@ 4.2 Ablation studies
We show ablation experiments in Figure 5. In the second column, we show results of our model where it is neither finetuned on the homography image nor uses mutual information guidance for sampling. Thus, the text embedding for the input image, followed by the diffusion UNet are finetuned and the diffusion model generates the aerial image by diffusion denoising, without any mutual information guidance. Many of the generated images either have low fidelity or have low correspondence to the aerial viewpoint. In the experiment in the third column, we add mutual information guidance to the model in column 2. We see higher fidelity (than column 2) of the generated images w.r.t. the input image. In the fourth column, we add the homography image finetuning step to the model in column 2, but do not use mutual information guidance at inference. The generated images, in many cases, are aerial, but have lower fidelity w.r.t. the input image. In the final column, we show results with our full model. The generated images achieve the best trade-off between the viewpoint being aerial and fidelity w.r.t. the input image, in comparison to all ablation experiments. We report a table with all quantitative metrics for the ablation experiments on AerialBooth-Syn and AerialBooth-Real in the supplementary pdf. A summary of the analysis is as follows:
  • Effect of $I_{H}$: To study the effect of $I_{H}$, we compare the A-CLIP, CLIPD, A-CLIPD scores for the full model vs model w/o $I_{H}$ and model w/o $I_{H}$ w/o $G_{MI}$ vs model w/o $G_{MI}$. The scores in all cases where $I_{H}$ is not present are lower. For instance, on AerialBooth-Real, the CLIP-D scores for the full model and model w/o $I_{H}$ are $0.0564$ and $0.0444$ respectively. Similarly, the CLIP-D scores for model w/o $I_{H}$ w/o $G_{MI}$ vs model w/o $G_{MI}$ are $0.0550$ and $0.0445$ respectively.
  • Effect of $G_{MI}$: To study the effect of $I_{H}$, we compare the SSCD, DINO, CLIPD, A-CLIPD scores for the full model vs model w/o $G_{MI}$ and model w/o $I_{H}$ w/o $G_{MI}$ vs model w/o $I_{H}$. The scores in all cases where $G_{MI}$ is not present are lower. For instance, on AerialBooth-Real, the SSCD scores for the full model and model w/o $G_{MI}$ are $0.3345$ and $0.3204$ respectively. Similarly, the CLIP-D scores for model w/o $I_{H}$ w/o $G_{MI}$ vs model w/o $I_{H}$ are $0.4013$ and $0.4066$ respectively.
Figure 6: AerialBooth can be extended to generate other text-controlled views as well.
Figure 7: We compare with latest related work on novel view synthesis: Zero-1-to-3[35] and Zero123++ [55]. Both methods use the pretrained stable diffusion model and the 3D objects dataset, Objaverse [10] for training. Our method uses just the pretrained stable diffusion model to generate better results for the task of aerial view synthesis from a single image.
@@ 4.3 Text controlled view synthesis: Other views
AerialBooth can be extended to generate other text-controlled views such as side view, bottom view and back view. For the results shown in Figure 19, we modify the target text $t_{T}$ to indicate different viewpoints, and retain the other finetuning/ inference details.
4.4 NVS Comparisons
We compare with latest work on stable diffusion based novel view synthesis from a single image in Figure 7. Zero-1-to-3[35] and Zero123++ [55], both, train on large amounts of multi-view and 3D data from Objaverse [10] contain $800K+$ 3D models; in addition to leveraging a pretrained stable diffusion model. In contrast, our method does not use any multi-view or 3D data for training and performs testtime finetuning to generate better results. However, while it is possible to explicitly control the camera angle and generate 3D objects in Zero-1-to-3[35] and Zero123++ [55], our method generates 2D images from arbitrary camera angles, within the realms of the text description of the viewpoint and consistent with the contents of the input image.
4.5 Comparison with other metrics for guidance.
We compare with two other metrics for diffusion guidance at inference: (i) L2 distance between the features of the generated image and the input image, (ii) a metric inspired by Wasserstein distance or Earth Mover’s distance, for which we compute the distance between the histograms of the probability distributions of the two images. Our mutual information guidance method is better at preserving the fidelity w.r.t the input image, as evidenced by higher SSCD scores. The SSCD score on AerialBooth-Real for Wasserstein guidance, $L2$ guidance, and mutual information guidance are $0.3181$, $0.3224$ and $0.3345$ respectively.
More results.
Please refer to the supplementary material for more analysis.
5 Conclusions, Limitations and Future

### v3

@@ 4 Experiments and Results
Figure 3: Compared to state-of-the-art text + exemplar image based methods, HawkI is able to generate images that are “more aerial”, while being consistent with the input image. The top three images are from the HawkI-Syn dataset, the bottom three images are from the HawkI-Real dataset.
Figure 4: HawkI achieves the best viewpoint-fidelity trade-off amongst prior work on text + exemplar image based aerial-view synthesis, on various quantitative metrics indicate of text-alignment (for viewpoint and a broad description of the scene) and image alignment (for fidelity w.r.t. input image).
@@
We collect a synthetic dataset, HawkI-Syn and a real dataset, HawkI-Real. Both datasets contain images across a wide variety of categories including indoor scenes, natural scenes, human actions, birds/ animals, animations, traffic scenes and architectures. HawkI-Syn contains $500$ images that were generated using Stable Diffusion XL [55]. To generate the text prompts for the generated images in HawkI-Syn, we used Large Languuage Models (LLMs) such as ChatGPT and Bard. HawkI-Real contains $139$ images downloaded from the Unsplash website, the text descriptions for these images were obtained using the BLIP-2 [40] model.
@@
We use the stable diffusion 2.1 model in all our experiments, ablations and comparisons. All our images (except for images in HawkI-Real) are at a resolution of $512\times 512$. With respect to $I_{S}$, we train the text embedding and the diffusion model for $1,000$ and $500$ iterations respectively, at a learning rate of $31e-$ 3 and $42e-$ 4 respectively. Using $1000$ iterations to optimize the text embedding ensures that the text embedding $e_{opt}$ at which $I_{S}$ is reconstructed is not too close to $e_{S}$, which would make it biased towards $I_{S}$ otherwise. Similarly, it is not too far from $e_{S}$ either, hence the text embedding space learns the characteristics of $I_{S}$. With respect to $I_{H}$, we train the text embedding and the diffusion UNet for $500$ and $250$ iterations respectively. We want $e_{H}$ to be in close vicinity of $e_{opt}$; we train the diffusion model for just $250$ iterations so that the model does not completely overfit to $I_{H}$. The role of $I_{H}$ is to create variance and provide pseudo supervision, it is not an accurate approximation of the aerial view. We set the hyperparameter for mutual information guidance at $51e-$ 5 or $61e-$ 6, the inference is run for $50$ steps.
Computational cost.
For each input image, HawkI takes $3.5$ minutes to perform test-time optimization on one NVIDIA A5000 GPU with 24 GB memory. The inference time is consistent with that of Stable Diffusion, about $7$ seconds to generate each sample with $50$ denoising steps. The computational cost is on par with text-based image personalization [65, 34] and text-based aerial-view synthesis [38] methods. The number of network parameters is also consistent across all these models, we use the Stable Diffusion v2.1 + LoRA backbone across methods.
@@
We follow prior work on text-based image editing/ personalization and text-based view synthesis to evaluate our method:
  • Viewpoint and text alignment: We use the text description ‘aerial view, ’ + $t_{S}$ and ‘aerial view’, along with the generated image, to compute the CLIP-Score [58] and the A-CLIP Score respectively. The former indicates alignment of the generated image with the detailed textual description of the image describing the contents along with the viewpoint; the latter focuses more on the viewpoint.
  • Image fidelity and 3D coherence: To evaluate the overall alignment of the contents of the generated aerial-view image with the input image, we compute the CLIP-I score [65] which measures the cosine similarity between the embeddings of the aerial-view image and the input image in the CLIP space. For a better indicator of the fidelity and 3D coherence between the two images, we also use the self-supervised similarity detection metrics, DINOv2 [8, 52] and SSCD [54].
  • Top-1 accuracy on a downstream UAV task.
  • Ground-truth comparison with images sampled from 3D models using CLIP, LPIPS, and DINO metrics.
Higher values are desired for each of these metrics. Viewpoint faithfulness and fidelity w.r.t input image are a direct result of the bias-variance trade-off of the model, and high values for both are desired. However, as noted by Blau et. al. [5], maximizing both is not straightforward; inevitably one of the factors will degrade in response to the improvement in the other. For each input image, we generate $5$ aerial images, with random noise initializations, and choose the image with the highest CLIP + SSCD score (since CLIP is an indicator of viewpoint + content alignment and SSCD score measures the fidelity w.r.t. the input image).
Figure 5: (Left figure.) Ablation experiments show that Inverse Perspective Mapping helps in the generation of images that are aerial, mutual information guidance helps in preserving the contents w.r.t. input image. (Right figure.) We compare with latest related work on novel view synthesis: Zero-1-to-3[48] and Zero123++ [70]. Both methods use the pretrained text-to-2D-image stable diffusion model along with the 800k 3D objects dataset, Objaverse [15], for training. Our method uses just the pretrained text-to-2D-image stable diffusion model to generate better results for the task of aerial view synthesis, guided by text and a single input image.
4.1 Comparisons against text + exemplar image based methods
@@
We show qualitative results in Figure 3. Our method is able to generate aerial views as per input image guidance across a diverse set of scenes. Our method generates results that are more aerial in viewpoint than DreamBooth and Imagic, while being largely consistent with the contents of the input image. Aerial Diffusion is unable to generate good quality images of scenes that have many objects. Our method is able to deal with complex scenes as well as modify the viewpoint.
@@
We show the quantitative results in Figure 7. Our method achieves a higher CLIP score than all prior work, indicating that it is able to generate an aerial view of the scene with contents dictated by the text better than prior work. The A-CLIP score achieved by our method is higher than that of DreamBooth and Imagic, indicating better conformance to the aerial viewpoint. Even though the A-CLIP score of HawkI is lower than than of Aerial Diffusion, Aerial Diffusion generates poor quality images for scenes with more than one object and also has diagonal artifacts in its generated images, as we observe from the other metrics and qualitative results, thus, offsetting its high A-CLIP Score.
@@
CLIP-I and self-supervised metrics such as SSCD and DINO are not viewpoint invariant. In many cases, since Imagic and DreamBooth generate views close to the input view, rather than aerial views, it is natural for them to have higher CLIP-I, SSCD and DINO Scores. Our method has a much higher CLIP-I, SSCD and DINO score than Aerial Diffusion, showing considerable improvement over prior work in retaining the fidelity and 3D consistency w.r.t. the input image, while modifying the viewpoint. In summary, our specialized aerial-view synthesis method achieves the best viewpoint-fidelity trade-off amongst all related prior work.
4.2 Comparisons against 3D based novel view synthesis (NVS) methods
We compare with state-of-the-art benchmark methods on stable diffusion based novel view synthesis from a single image in Figure 5. Zero-1-to-3[48] and Zero123++ [70], both, train on large amounts of multi-view and 3D data from Objaverse [15] contain $800K+$ 3D models; in addition to leveraging a pretrained text-to-2Dimage stable diffusion model. In contrast, our method does not use any multi-view or 3D information and is capable of generating better results in multiple cases. Another task-level difference between our method and prior work on NVS is that the latter aim to explicitly control the camera angle and generate 3D objects in Zero-1-to-3[48], the camera-angle generated by our method is arbitrary within the realms of the text control. The CLIP scores on HawkI-Syn for Zero123++ and HawkI are $0.3071$ and $0.3115$ respectively, the DINO scores are $0.4341$ and $0.4466$ respectively. On HawkI-Real, the CLIP score for Zero123++ and HawkI are $0.2908$ and $0.3077$ respectively, the DINO scores are $0.3916$ and $0.3956$ respectively. Our aerial-view synthesis method, even without any 3D/ multi-view information and large dataset training, is better than or comparable to 3D-based NVS methods.
4.3 Downstream application.
We performed a proof-of-concept experiment using HawkI to generate synthetic images for key-frames of scene-based human actions in the UAV Human [41] dataset. By computing the L2 distance between self-supervised features of UAV Human video frames and synthetic images on actions: ‘smoking’, ‘pushing someone’, ‘high five’ and ‘walking’, we achieved a $32.14\%$ accuracy in zero-shot action recognition, representing a $7.14\%$ (absolute) improvement.
4.4 Ground-truth comparison
We obtained 3D models and text descriptions from Dreamfusion [56] to extract the front-view and top-view (ground-truth or GT). We evaluated Zero123++ and HawkI using CLIP Scores (higher the better), and the numbers were $0.2991$, $0.3316$, respectively. The DINO score (higher the better), which measures self-supervised similarity between the generated images and the GT, are $0.3912$, $0.3710$ for Zero123++ and HawkI respectively. The LPIPS scores (lower the better) are $0.5801$, $0.6373$ for Zero123++ and HawkI respectively - our method generates images that are higher in elevation due to ‘aerial view’ being the text-control. Overall, our 3D-free method, built on stable diffusion, is comparable to 3D methods such as Zero123++ which uses stable diffusion + 800k+ 3D objects.
@@ 4.5 Ablation studies
We show ablation experiments in Figure 5. In the second column, we show results of our model where it is neither finetuned on the homography image nor uses mutual information guidance for sampling. Thus, the text embedding for the input image, followed by the diffusion UNet are finetuned and the diffusion model generates the aerial image by diffusion denoising, without any mutual information guidance. Many of the generated images either have low fidelity or have low correspondence to the aerial viewpoint. In the experiment in the third column, we add mutual information guidance to the model in column 2. We see higher fidelity (than column 2) of the generated images w.r.t. the input image. In the fourth column, we add the homography image finetuning step to the model in column 2, but do not use mutual information guidance at inference. The generated images, in many cases, are aerial, but have lower fidelity w.r.t. the input image. In the final column, we show results with our full model. The generated images achieve the best trade-off between the viewpoint being aerial and fidelity w.r.t. the input image, in comparison to all ablation experiments. Quantitative analysis:
  • Effect of $I_{H}$: To study the effect of $I_{H}$, we compare the CLIP and A-CLIP scores for the full model vs full model w/o $I_{H}$. The scores in all cases where $I_{H}$ is not present are lower, indicating lower consistency with the viewpoint being ‘aerial’. For instance, on HawkI-Real, the CLIP scores for the full model and full model w/o $G_{MI}$ are $0.3077$ and $0.3040$ respectively. The A-CLIP scores for the full model and full model w/o $I_{H}$ are $0.1887$ and $0.1842$ respectively.
  • Effect of $G_{MI}$: To study the effect of $I_{H}$, we compare the SSCD and DINO scores for the full model vs model w/o $G_{MI}$. The scores in all cases where $G_{MI}$ is not present are lower. For instance, on HawkI-Real, the SSCD scores for the full model and model w/o $G_{MI}$ are $0.3314$ and $0.3204$ respectively. The DINO scores for the full model and model w/o $G_{MI}$ are $0.3956$ and $0.39$ respectively.
Figure 6: HawkI can be extended to generate other text-controlled views as well.
4.6 Comparison with other metrics for guidance.
We compare with two other metrics for diffusion guidance at inference: (i) L2 distance between the features of the generated image and the input image, (ii) a metric inspired by Wasserstein distance or Earth Mover’s distance, for which we compute the distance between the histograms of the probability distributions of the two images. Our mutual information guidance method is better at preserving the fidelity w.r.t the input image, as evidenced by higher SSCD scores. The SSCD score on HawkI-Real for Wasserstein guidance, $L2$ guidance, and mutual information guidance are $0.3137$, $0.3204$ and $0.3314$ respectively. The DINO score on HawkI-Real for Wasserstein guidance, $L2$ guidance, and mutual information guidance are $0.3847$, $0.3858$ and $0.3956$ respectively.
@@ 4.7 Text controlled view synthesis: Other views
HawkI can be extended to generate other text-controlled views such as side view, bottom view and back view. For the results shown in Figure 19, we modify the target text $t_{T}$ to indicate different viewpoints, and retain the other finetuning/ inference details.
4.8 3D-free HawkI + 3D priors?
Our 3D-free approach complements 3D-based methods [70, 42]. While 3D data collection and large-scale training are expensive and unsustainable, front-view 2D images are more readily available[35]. Hence, it is beneficial to solve the fundamental problems associated with the task in a data-efficient (or 3D-free) manner. Our goal is to push the frontiers of 3D-free aerial-view generation from a single image, achieving results comparable to methods [70] using stable diffusion + 800k 3D objects. That being said, HawkI be combined with 3D approaches, to multiply the benefits of 3D and 3D-free approaches. To validate this, we replaced the homography prior with the image from the 3D-based Zero123++ [70] approach and evaluated our HawkI model on images from HawkI-Syn. The CLIP scores for Zero123++, HawkI, and “HawkI with Zero123++ prior” are $33.17$, $33.17$, and $33.81$, respectively. The DINO scores for Zero123++, HawkI, and “HawkI with Zero123++ prior” are $0.3613$, $0.3977$, and $0.4612$, respectively, thus demonstrating our claim!. Also, using 3D priors with HawkI allows finer camera control.
More results. Please refer to the supplementary material for (i) more qualitative comparisons with text + exemplar image and 3D based NVS methods, (ii) qualitative comparisons with other guidance metrics, (iii) detailed quantitative results for ablations, (iv) comparison of IPM with data augmentation, (v) more results on other text-controlled views, (vi) qualitative comparisons with warping + outpainting (scene extrapolation) and ControlNet variations.
5 Conclusions, Limitations and Future Work